OpenStudy (jazilove):

PLEASE HELP WILL MEDAL If I started with 15 grams of water and it decomposed into hydrogen and oxygen, how many grams of hydrogen and oxygen would I have?

OpenStudy (sapphiremoon):

This is stoichiometry. First, you'll want to find out how many moles of water you have: \[15g H_{2}O \times \frac{ 1 mol }{ 18.02 g } = 0.8324 mol H_{2}O\] (sorry, the equations thing doesn't do spaces...) Then we find out how many mol of H and O we have: \[0.8324 mol H_{2}O \times \frac{ 2 mol H }{ 1 mol H_{2}O } = 1.664 mol H\] \[0.8324 mol H_{2}O \times \frac{ 1 mol O }{ 1 mol H_{2}O } = 0.8324 mol O\] Now we convert the mols of each back into grams. \[1.664 mol H \times \frac{ 1.01 g H }{ mol H} = 1.6806 g H\] \[0.8324 mol O \times \frac{ 16.0 g O }{ mol O} = 13.318 g O\] Which round to 1.7g H and 13 g O respectively based on significant figures. Does this make sense? Any questions?

OpenStudy (jazilove):

I am so sorry I had to step away, but this helps so much thank you !

OpenStudy (sapphiremoon):

All right. You're welcome!